Circular No. 8435

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2004fy NEAR MCG +15-1-10 Further to IAUC 8431, T. Puckett and B. Kerns report the discovery of

an apparent supernova (mag 17.0) on unfiltered CCD frames taken with the 0.35-m automated supernova patrol telescope on Nov. 10.46 and 11.06 UT. The new object is located at α = 11^{h}41^{m}14^{s}.96, δ = +89^{o}05′10′′.6 (equinox 2000.0), which is 9′′.2 east and 5′′.7 south of the center of MCG +15-1-10. Nothing is visible at this location on Palomar Sky Survey images taken on 1988 Jan. 28 and 1998 Apr. 22 (limiting mag ∼ 21.0).

V4743 SAGITTARII M. Orio, Italian National Institute for Astrophysics, Turin, and Univer-

sity of Wisconsin; and E. Tepedelenlioglu, University of Wisconsin, report that V4743 Sgr was observed with the XMM-Newton satellite beginning on Sept. 30.783 UT for ∼ 6.2 hr (two years after the optical outburst). The EPIC-pn background-corrected count rate was 0.347 ± 0.007 counts/s in the range 0.3–10 keV in timing mode, while count rates of 0.101 ± 0.002 and 0.107 ± 0.002 counts/s were measured with EPIC-MOS-1 and EPIC-MOS-2, respectively, in the range 0.2–10 keV. The average absorbed flux in the range 0.2–10 keV is ∼ 1.3 × 10^{−12} erg cm^{−2} s^{−1}, a factor of 1000 lower than the flux measured in March and April 2003 (IAUC 8107, 8131). The nova still appears as a supersoft x-ray source, and deep absorption features due to N VII and N VI are still detected, but there is also a hard component, possibly due to the ejecta. Preliminary results from the lightcurve analysis show again several peaks in the power spectrum, the highest of them at the frequency corresponding to a 1365-s period.

SUPERNOVA 2004et IN NGC 6946 R. J. Beswick, T. W. B. Muxlow, M. K. Argo, and A. Pedlar, Jodrell

Bank Observatory; and J. M. Marcaide, University of Valencia, report the radio detection of SN 2004et at 4.994 GHz, using a subset of the MERLIN array. A compact radio source at α = 20^{h}35^{m}25^{s}.360, δ = +60^{o}07′17′′.74 (equinox 2000.0, ± 0′′.04 in each coordinate) is consistent with the radio position reported by Stockdale et al. (IAUC 8415). MERLIN 4.994-GHz radio-flux measurements for SN 2004et are: Sept. 30–Oct. 13, 0.4 ± 0.3 mJy; Oct. 15–17, 0.61 ± 0.4; Oct. 18–20, 1.55 ± 0.3; Oct. 25–27, 1.31 ± 0.4; Oct. 27–28, 1.34 ± 0.4; Oct. 29–30, 1.55 ± 0.3; Oct. 31–Nov. 1, 1.89 ± 0.3, Nov. 2, 2.39 ± 0.3. MERLIN and VLA radio monitoring is continuing.

2004 November 11  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2004 CBAT  Daniel W. E. Green